OpenStudy (anonymous):

log 8^(3x - 1) = (1/5)^x need help again!

zepdrix (zepdrix):

I'll have something for you in a sec :) Sorry I have to type it in a text pad. OpenStudy keeps crashing when I try to write in this box.

zepdrix (zepdrix):

\[\Large 8^{(3x - 1)} = \left(\frac{1}{5}\right)^x\]Taking the Natural Log of both sides gives us,\[\Large \ln\left(8^{(3x - 1)}\right)=\ln\left[\left(\frac{1}{5}\right)^x\right]\]Applying rules of logs allows us to write it like this,\[\Large (3x-1)\ln8=x\ln\left[\frac{1}{5}\right]\]

zepdrix (zepdrix):

Lemme know if you're confused by any of that so far :o

OpenStudy (anonymous):

so far i understand. let me try to solve it from here.

zepdrix (zepdrix):

Since we could have used any log to solve this, (not just the natural log), final answers will vary way too much. So it's probably best to get a decimal approximation :) Don't leave it in exact form.

OpenStudy (anonymous):

but i don't know how to move forward from this ... (3x-1)In8 = -xIn5 ;I ... am I right so far ?

zepdrix (zepdrix):

The next step is a little tricky. Requires some clever algebra.

zepdrix (zepdrix):

Distribute the ln8 to each term inside the brackets.

zepdrix (zepdrix):

Mmm I like the step you did with the 1/5. It's not necessary, but it makes that log a little easier to put into the calculator.

zepdrix (zepdrix):

\[\Large 3x\ln8-\ln8=-x\ln5\]Understand how we did that?

OpenStudy (anonymous):

ok let me try. Sorry I was away for a while, to save my cat....

zepdrix (zepdrix):

oh no! Mr kitty!! \:o/

OpenStudy (anonymous):

i feel very stupid... =,( i can't see the steps to solution. like this? 3xIn8 + xIn5 = In8.

zepdrix (zepdrix):

Good :) Now on the left side, let's factor x out of each term.

OpenStudy (anonymous):

x(3In8+In5) = In8?

zepdrix (zepdrix):

K looks good. Now divide both sides by the coefficient on x. That big messy thing, move it to the other side.

OpenStudy (anonymous):

can I do like this? x = In8/(3In8+In5) x = 2.08/[3(2.08) + 1.61)]

zepdrix (zepdrix):

yesss very good!

OpenStudy (anonymous):

wao! finally haha. thank you so much for your help! I need to do more to understand the rules of log....

zepdrix (zepdrix):

Those weird algebra tricks are so annoying! :) lol

OpenStudy (anonymous):

agree! u are so smart to see the tricks!